OpenStudy (anonymous):

solve by elimination: 4x=10+2y -6/5x+y=-1/5

OpenStudy (anonymous):

solve by elimination: 4x=10+2y y=2x-5

OpenStudy (anonymous):

5xy-6/5x=-1/5

OpenStudy (experimentx):

first fix your equation so that it would be easier 4x - 2y =10 -6/5x + y=-1/5 4x - 2y =10 2(-6/5x + y)=2(-1/5) eleminate y, 4x+2(-6/5x) = 10+(-1/5) rest ... i hope you can do

OpenStudy (anonymous):

thanks e